OpenStudy (anonymous):

How do you find the sum of the series of 3^(k-1)/4^(3k+1) ; k = 0 to infinity? (infinite series)

OpenStudy (anonymous):

You just have to do some manipulation on the series to see that it's actually geometric, and then all you need to do is use the sum of a geometric series formula. So...

OpenStudy (anonymous):

\[\sum_{k=0}^{\infty}\frac{3^{k-1}}{4^{3k+1}}=\sum_{k=0}^{\infty}\frac{3^k.3^{-1}}{(4^3)^k.4}=\frac{1}{12}\sum_{k=0}^{k=\infty}(\frac{3}{64})^k\]

OpenStudy (anonymous):

The geometric series has first term a=1 and common ratio r=3/64. The sum is then,\[\frac{a}{1-r}=\frac{1}{1-3/64}=\frac{64}{61}\]

OpenStudy (anonymous):

You have to now multiply this answer by 1/12 to get your sum:\[\sum_{k=0}^{\infty}\frac{3^{k-1}}{4^{3k+1}}=\frac{16}{183}\]

OpenStudy (anonymous):

Please fan me :) I punch out answers and no one bothers to reward! Cheers.

OpenStudy (anonymous):

Wow thank you and how do I fan you? I am new to this site

OpenStudy (anonymous):

There should be a [thumbs up] icon next to my name and the words, "Become a fan" underlined as a link. You just click on that...it's blue...

OpenStudy (anonymous):

oh lol I had refresh it to see the 'become a fan' button but anyway thank you for your help!

OpenStudy (anonymous):

cheers...no worries.